4^{th} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS

Krynica 29th April – 04th May 2002

http://im0.p.lodz.pl/konferencje/krynica2002

Organized by

– Institute of Mathematics of the Technical University of Lodz Address: Al. Politechniki 11, 90-924 Łódź Jan Kubarski (chairman), e-mail: kubarski@ck-sg.p.lodz.pl

– Institute of Mathematics of the Jagiellonian University, Cracow Address: ul. Wł. Reymonta 4, 30-059 Kraków Robert Wolak, e-mail: wolak@im.uj.edu.pl

– Faculty of Applied Mathematics of the University of Mining and Metallurgy , Cracow Address: al. Mickiewicza 30, 30-059 Kraków, Tomasz Rybicki, e-mail: tomasz@uci.agh.edu.pl

THIRD ANNOUNCEMENT

Dear Colleague

28 IV (the Sunday) is planed as the day of arriving. The University coach will leave the Institute of Mathematics of the Jagiellonnian University at 3.30 p.m. : Wl. Reymonta Str. 4, phone: 48-12-6339781. Our first common meal will be the dinner at 7.30 p.m. at ”Baśka” hotel in Krynica (phone: 018 4715888) which is situated on a hill overlooking the railway station in Krynica.

Kraków can be reach easily from Warsaw by fast trains which leave Warszawa Centralna railway station at 6.15, 6.50 7.35, 8.35, 11.35, 12.35, 15.35, 16.35, 17.35, 19.35 . The journey, without any stop, takes about 2.40 h, unfortunately the reservation is obligatory. From Kraków Główny you should take tram no 15 and descend at the fifth stop at the corner of Oleandry Str. At the end of this street (about 200 m) there is the building of the Institute. The ticket can be bought from the driver and costs 2.70 PLN, in kiosks 2.20 PLN.

There are two direct trains from Warsaw to Krynica: dep. Warszawa Centralna 6:15, arr. Krynica 13:12 dep. Warszawa Zachodnia 21:30, arr. Krynica 7:09.

In case of any problem please contact Robert Wolak at wolak@im.uj.edu.pl or +48601546642@text.plusgsm.pl (e-mail address of the mobile phone) so that we can plan your journey to Krynica. You can also contact Robert Wolak on his mobile phone: 0048601546642 from outside Poland or 0601546642 in Poland.

If you have not sent yet the full registration or the title of your lecture or the abstract please do it as quickly as possible.

The conference is included in the world Atlas of conferences http://atlas-conferences.com/ Therefore our conference has been set up to use the services of Atlas Mathe-

matical Conference Abstracts. Please be so kind and submit also your abstract at

http://at.yorku.ca/cgi-bin/amca/submit/cajc-01 and to view submitted abstracts at http://at.yorku.ca/cgi-bin/amca/cajc-01 We plan to publish the proceedings in the Journal ”Universitatis Iagellonicae

Acta Mathematica”. Details will be given at the Conference. The deadline for submitted our papers is the 1 October 2002.

Best regards, organizers:

Jan Kubarski, Robert Wolak, Tomasz Rybicki